# arXiv:2603.00301v1[math.NT] 27 Feb 2026

### 27 Feb 2026 v1

@@ 4.4. Cesàro spacelike theorem
Reproducibility. The bilinear form $h(T)=\sum_{j,k}b_{j}b_{k}\,F(\gamma_{j}^{\prime},\gamma_{k}^{\prime},T)$ was evaluated using python-flint (ARB interval arithmetic) at 512-bit working precision. For each $d$, the function $h(T)$ was decomposed into a constant plus $N(N+1)/2$ sinc terms via the identity $h(T)=C+\sum_{m}A_{m}\operatorname{sinc}(\omega_{m}T)$, where the frequencies $\omega_{m}$ are all pairwise sums and differences $\gamma_{j}^{\prime}\pm\gamma_{k}^{\prime}$ and the coefficients $A_{m}$ are the corresponding weight products. The grid scan used spacing $5\times 10^{-4}$ over $T\in[0.01,50]$; the Lipschitz bound $|h^{\prime}(T)|\leq\frac{1}{2}\sum_{j,k}b_{j}b_{k}\max(\gamma_{j}^{\prime},\gamma_{k}^{\prime})$ uses the elementary estimate $|(\operatorname{sinc})^{\prime}(x)|\leq\frac{1}{2}$ for all $x\geq 0$. Every grid evaluation produces a certified interval with radius below $10^{-150}$; the certified minimum is the smallest lower endpoint minus the Lipschitz discretization error $L\Delta$. The large-$T$ tail is handled by the analytic bound from Case 1, with all constants ($c_{0}$, $W_{2}$, $\mathcal{I}$, $T_{\mathrm{crit}}$) computed in ARB. The worst-case five-zero configuration for $d\geq 14$ used equally spaced frequencies $\gamma_{k}^{\prime}\in\{6.0,7.25,8.5,9.75,11.0\}$ with $b_{k}=2/(\tfrac{1}{4}+{\gamma_{k}^{\prime}}^{2})$, minimized over the same grid.
@@ 5. Besicovitch Norms of Spectral Sums
Sections 5 through 7 establish the density bounds; the reader primarily interested in the unconditional results may proceed directly to Theorem 8.5.
The functions $S_{\zeta}$ and $S_{L}$ live in the Besicovitch Hilbert space $B^{2}$ with inner product
@@ A.1. Software and rigorous arithmetic
Working precisions: 512-bit ($\approx 154$ decimal digits) for $S_{\zeta}^{*}$ certification, spacelike verification (Table 2), and null crossing computation (ARB); 500-bit for auxiliary $L$-function zeros (ARB); 1500-bit ($\approx 451$ decimal digits) for all $\chi_{5}$ zero refinement and certification (ARB); 256-bit for all Bessel product integrals including density, resonance, and decay estimates (ARB); and 80-digit decimal precision for PSLQ searches (mpmath).
@@ A.2. $L$-function evaluation and zero-finding
Stage 1: Sign-change scan. Evaluate $Z_{\chi}(t)$ on a uniform grid with step $0.25$ over $0,T_{\max}]$, where $T_{\max}$ is chosen to exceed the expected height of the last zero sought. The zero-counting formula $N(T,\chi_{d})\sim\frac{T}{\pi}\log\frac{qT}{2\pi e}$ provides a conservative estimate; for example, $200$ zeros of $L(s,\chi_{5})$ require $T_{\max}\approx 290$ (the asymptotic formula gives $N(290,\chi_{5})\approx 205$). Low precision (128-bit ARB) suffices since only the sign of $Z_{\chi}$ is required. The scan produces $\lceil T_{\max}/0.25\rceil$ evaluations; for $T_{\max}=290$ this is $1160$ evaluations, each taking a fraction of a second at 128-bit precision.
Stage 2: Bisection. For each consecutive pair $t_{a},t_{b}$ with $Z_{\chi}(t_{a})\,Z_{\chi}(t_{b})<0$, bisect on $Z_{\chi}$ to isolate the root. At 512-bit ARB precision, $100$ bisection steps produce $\sim$30 decimal digits from a bracket of width $0.25$, yielding seeds accurate to at least 20 decimal places. Certification at this stage: accept $\gamma^{\prime}$ when $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|<10^{-30}$. (Note: Stages 1 and 2 need be performed only once per character; the resulting 20-digit seeds are tabulated in Appendix C and may be reused indefinitely as input to Stage 3.)
Stage 3: Newton refinement. Refine each seed by Newton iteration applied to $Z_{\chi}(t)$. The derivative is approximated by the finite difference $(Z_{\chi}(t+h)-Z_{\chi}(t))/h$ with $h=10^{-50}$; the midpoint of the resulting ARB ball is extracted after each Newton step (see Remark A.1 below). At 1500-bit precision, $8$ Newton iterations from a 20-digit seed converge to full working precision ($\approx 451$ digits), with certified bounds on $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|$ typically in the range $10^{-358}$$10^{-400}$. The convergence is quadratic: each iteration approximately doubles the number of correct digits, so $20\to 40\to 80\to 160\to 320\to 451$ digits in five to six effective steps.
Certification. A zero ordinate $\gamma^{\prime}$ is accepted when $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|<\text{threshold}$, with the threshold depending on application (typically $10^{-30}$ for 20-digit values, $10^{-68}$ for 70-digit values). The certified bound is computed by evaluating $L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})$ in ARB ball arithmetic and extracting the upper endpoint of the absolute value ball; this is a rigorous upper bound as a mathematical theorem. Readers who already possess 20-digit seeds (for example, from the tables in Appendix C) may skip directly to Stage 3.
Precomputation. The constants $\log q$, $\log(q/\pi)$, and the Hurwitz parameters $a/q$ for each nonzero character value should be computed once and stored as arb objects at the working precision. This avoids redundant string-to-ball conversions in the inner loop and reduces the per-evaluation cost by approximately $30\%$.
@@
(3) The finite-difference step $h$ must be appropriate for the current precision. The step $h=10^{-50}$ is suitable for 1500-bit ($\approx 451$-digit) working precision: it is large enough to avoid catastrophic cancellation (the difference Z(t+h)Z(t)Z(t+h)-Z(t) retains $\approx 400$ significant digits) and small enough to make the $O(h)$ truncation error negligible relative to the quadratic convergence. A step that is too small (for example, $h=10^{-200}$ at 512-bit precision) produces a derivative estimate dominated by rounding error, causing Newton iteration to stall.
@@ A.4. Verification of spacelike constraint
22 88 33 $0.13376$ $Z_{\chi}$ computation
55 55 22 $0.06563$ $Z_{\chi}$ computation
$q\geq 12$ $\geq 12$ $\geq 4$ $0.06598$ BMOR bound
@@ A.6. Bessel product integrals
For the density integral (14): $f_{S_{L}}(0)=\frac{1}{\pi}\int_{0}^{T}\prod_{k=1}^{20}J_{0}(b_{k}t)\,dt$ with $T=2000$. The Lorentzian weights $b_{k}=2/(\tfrac{1}{4}+{\gamma_{k}^{\prime}}^{2})$ are computed in ARB from the 70-digit certified zeros of Appendix B. The integrand is entire (products of $J_{0}$), so no branch cut handling is required. ARB quadrature at 256-bit precision yields the certified result:

### 03 Mar 2026 v2

@@ 4.4. Cesàro spacelike theorem
Reproducibility. The bilinear form $h(T)=\sum_{j,k}b_{j}b_{k}\,F(\gamma_{j}^{\prime},\gamma_{k}^{\prime},T)$ was evaluated using python-flint (ARB interval arithmetic) at 512-bit working precision. For each $d$, the function $h(T)$ was decomposed into a constant plus $N(N+1)/2$ sinc terms via the identity $h(T)=C+\sum_{m}A_{m}\operatorname{sinc}(\omega_{m}T)$, where the frequencies $\omega_{m}$ are all pairwise sums and differences $\gamma_{j}^{\prime}\pm\gamma_{k}^{\prime}$ and the coefficients $A_{m}$ are the corresponding weight products. The grid scan used spacing $5\times 10^{-4}$ over $T\in[0.01,50]$; the Lipschitz bound $|h^{\prime}(T)|\leq\frac{1}{2}\sum_{j,k}b_{j}b_{k}\max(\gamma_{j}^{\prime},\gamma_{k}^{\prime})$ uses the elementary estimate $|(\operatorname{sinc})^{\prime}(x)|\leq\frac{1}{2}$ for all $x\geq 0$. Every grid evaluation produces a certified interval with radius below $10^{-150}$; the certified minimum is the smallest lower endpoint minus the Lipschitz discretization error $L\Delta$. The large-$T$ tail is handled by the analytic bound from Case 1, with all constants ($c_{0}$, $W_{2}$, $\mathcal{I}$, $T_{\mathrm{crit}}$) computed in ARB. The worst-case five-zero configuration for $d\geq 14$ used equally spaced $\gamma^{\prime}_{\mathrm{unk}}$ frequencies $\gamma_{k}^{\prime}\in\{6.0,7.25,8.5,9.75,11.0\}$ with $b_{k}=w^{-}(\gamma_{k}^{\prime})=2/(1+{\gamma_{k}^{\prime}}^{2})$, minimized over the same grid.
@@ 5. Besicovitch Norms of Spectral Sums
The next three sections establish that the truncated spectral sum $S_{L}^{(M)}$ has a Cesàro distribution with finite density at the origin, and use this to prove the effective density bound (Theorem 8.5). The reader primarily interested in the unconditional density bound may proceed directly to that theorem; the present section and the two that follow supply its proof.
Three obstacles must be overcome. First, the Cesàro characteristic function of $S_{L}^{(M)}$ is a product of Bessel functions $J_{0}(b_{k}t)$, and the integral $(1/\pi)\int_{0}^{\infty}\prod|J_{0}(b_{k}t)|\,dt$ must be shown to converge; this is the content of the Bessel decay estimates in the present section, and it succeeds because the product decays as $t^{-M/2}$ for $M\geq 3$. Second, the characteristic function is not simply the Bessel product: integer relations among zero ordinates (resonances) contribute correction terms indexed by a lattice $\Lambda_{M}$, and each such term must be shown to decay exponentially in the order of the relation. The subcritical and transition Bessel estimates of §6.3, combined with the lattice counting of §6.4, control these corrections unconditionally at every finite truncation level $M$ (Theorem 6.9). Third, passing from finite $M$ to the full infinite series requires that the resonance corrections remain summable as $M\to\infty$; this is the telescoping extension of Section 7, which exploits a self-referential structure whereby each new zero’s smallness suppresses its own resonance contribution.
We begin by placing the spectral sums in the Besicovitch Hilbert space, which provides the natural inner product for Cesàro-distributed almost periodic functions. The functions $S_{\zeta}$ and $S_{L}$ live in the Besicovitch Hilbert space $B^{2}$ with inner product
@@ 7.3. The major-arc/minor-arc decomposition
Remark 7.9 (Structural nature of the finite-to-infinite obstruction).
It is natural to ask whether the passage from verified resonance absence at finite $M$ to the limiting density $f_{S_{L}}(0)<\infty$ can be achieved by purely approximation-theoretic means, bypassing the hypothesis $d_{\infty}<\infty$. Since $S_{L}^{(M)}\to S_{L}$ in $B^{2}$, the value distributions converge weakly (by Cauchy–Schwarz applied to characteristic functions and Lévy continuity). However, weak convergence does not imply pointwise convergence of densities. The Fourier inversion formula $f^{(M)}(0)=(2\pi)^{-1}\int\prod_{k=1}^{M}J_{0}(b_{k}t)\,dt$ requires the product factorization of the characteristic function, which in turn requires $\mathbb{Q}$-linear independence of the frequencies $\gamma_{1}^{\prime},\ldots,\gamma_{M}^{\prime}$. The dominated convergence argument that yields $f^{(M)}(0)\to f^{(\infty)}(0)$ relies on the monotone bound $|\varphi^{(M+1)}(t)|\leq|\varphi^{(M)}(t)|$, which holds only under the product structure. Without independence, the characteristic function need not factor, the monotone domination fails, and no alternative $L^{1}$ dominator is available. The obstruction is therefore structural, not computational: every Fourier-analytic path to $f_{S_{L}}(0)<\infty$ passes through the independence of zero ordinates.
@@ A.1. Software and rigorous arithmetic
Working precisions: 512-bit ($\approx 154$ decimal digits) for $S_{\zeta}^{*}$ certification, spacelike verification (Table 2), and null crossing computation (ARB); 500-bit for the auxiliary $L$-function zeros tabulated in Appendices C and E (20 decimal places, computed with an earlier version of the zero-finding pipeline); 1500-bit ($\approx 451$ decimal digits) for all $\chi_{5}$ zero refinement and certification (ARB); 128-bit for the sign-change scan and bracket filtering phases described in §A.2, used in the extended 1000-zero computations of Appendix F; 256-bit for all Bessel product integrals including density, resonance, and decay estimates (ARB); and 80-digit decimal precision for PSLQ searches (mpmath).
@@ A.2. $L$-function evaluation and zero-finding
Phases 1 and 2: Interleaved sign-change scan and bracket filtering. Evaluate $Z_{\chi}(t)$ on a uniform grid with step $0.05$ over $0,T_{\max}]$, where $T_{\max}$ is chosen so that the zero-counting formula $N(T,\chi_{d})\sim\frac{T}{\pi}\log\frac{qT}{2\pi e}$ exceeds the number of zeros sought by a safety margin; for example, $200$ zeros of $L(s,\chi_{5})$ require $T_{\max}\approx 290$. All arithmetic in this phase uses 128-bit ARB precision. At each sign change $Z_{\chi}(t_{a})\,Z_{\chi}(t_{b})<0$, $20$ bisection steps are performed immediately (interleaved with the grid scan), reducing the bracket width from $0.05$ to $0.05\cdot 2^{-20}\approx 5\times 10^{-8}$ and producing a seed accurate to approximately six decimal digits. The bisection simultaneously filters spurious sign changes arising from Hardy-phase oscillations: at the bisected midpoint, $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|$ is evaluated and the bracket is accepted as a genuine zero only if $|L|<10^{-4}$; brackets with $|L|>10^{-3}$ are discarded. This interleaved design avoids storing all sign changes before filtering, so scanning stops as soon as the required number of genuine seeds has been collected.
Phase 3: Newton refinement. Refine each seed by Newton iteration applied to $Z_{\chi}(t)$. The derivative is approximated by the finite difference $(Z_{\chi}(t+h)-Z_{\chi}(t))/h$; the midpoint of the resulting ARB ball is extracted after each Newton step (see Remark A.1 below). Two modes are used depending on the target precision:
  • Standard mode (128-bit, $h=10^{-8}$): from a six-digit seed, quadratic convergence reaches $\sim$20 correct digits in two to three iterations. Certification threshold: $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|<10^{-18}$.
  • High-precision mode (1500-bit, $h=10^{-50}$): from a six-digit seed, seven to eight iterations reach full working precision ($\approx 451$ digits), with certified bounds on $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|$ typically in the range $10^{-358}$$10^{-450}$. Certification threshold: $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|<10^{-68}$.
In both modes, the convergence is quadratic: each iteration approximately doubles the number of correct digits.
Certification. A zero ordinate $\gamma^{\prime}$ is accepted when $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})|<\text{threshold}$, with the threshold depending on the target precision ($10^{-18}$ for 20-digit values, $10^{-68}$ for 70-digit values). The certified bound is computed by evaluating $L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{d})$ in ARB ball arithmetic and extracting the upper endpoint of the absolute value ball; this is a rigorous upper bound as a mathematical theorem.
Precomputation. The constants $\log q$, $\log(q/\pi)$, and the Hurwitz parameters $a/q$ for each nonzero character value are computed once and stored as arb objects at the working precision. This avoids redundant string-to-ball conversions in the inner loop and reduces the per-evaluation cost by approximately $30\%$.
@@
(3) The finite-difference step $h$ must be appropriate for the current precision. The step $h=10^{-50}$ is suitable for 1500-bit ($\approx 451$-digit) working precision: it is large enough to avoid catastrophic cancellation (the difference Z(t+h)Z(t)Z(t+h)-Z(t) retains $\approx 400$ significant digits) and small enough to make the $O(h)$ truncation error negligible relative to the quadratic convergence. For 128-bit standard mode, $h=10^{-8}$ provides the analogous balance at lower precision. A step that is too small for the working precision (for example, $h=10^{-200}$ at 512-bit) produces a derivative estimate dominated by rounding error, causing Newton iteration to stall.
@@ A.4. Verification of spacelike constraint
22 88 33 $0.13376$ $Z_{\chi}$: $\gamma^{\prime}_{\mathrm{knw}}$, $w=2/(\tfrac{1}{4}+\gamma^{\prime 2})$
55 55 22 $0.06563$ $Z_{\chi}$: $\gamma^{\prime}_{\mathrm{knw}}$, $w=2/(\tfrac{1}{4}+\gamma^{\prime 2})$
$q\geq 12$ $\geq 12$ $\geq 4$ $0.06557$ BMOR: $\gamma^{\prime}_{\mathrm{unk}}$, $w^{-}=2/(1+\gamma^{\prime 2})$
@@ A.6. Bessel product integrals
For the density integral (15): $f_{S_{L}}(0)=\frac{1}{\pi}\int_{0}^{T}\prod_{k=1}^{20}J_{0}(b_{k}t)\,dt$ with $T=2000$. The Lorentzian weights $b_{k}=2/(\tfrac{1}{4}+{\gamma_{k}^{\prime}}^{2})$ are computed in ARB from the 70-digit certified $\gamma^{\prime}_{\mathrm{knw}}$ zeros of Appendix B. The integrand is entire (products of $J_{0}$), so no branch cut handling is required. ARB quadrature at 256-bit precision yields the certified result:
@@ Appendix E ARB Certification of Zero Data
Appendix F Tables of Zeros of Quadratic Dirichlet $L$-Functions of Small Conductor
Due to its length (approximately 181 pages), this appendix is archived separately on Zenodo:
Appendix F: Tables of Zeros of Quadratic Dirichlet $L$-Functions of Small Conductor.
Available at https://doi.org/10.5281/zenodo.18783098.
The companion document tabulates the first 1000 zeros of $L(s,\chi_{d})$ to 70 decimal places for each of the eight quadratic characters $\chi_{2}$, $\chi_{3}$, $\chi_{5}$, $\chi_{6}$, $\chi_{7}$, $\chi_{10}$, $\chi_{11}$, and $\chi_{13}$ (8000 zeros in total). All zeros were computed using the methodology of §A.2 at 1500-bit working precision and rigorously certified using ARB interval arithmetic (§A.1). Certified upper bounds on $|L(\tfrac{1}{2}+i\gamma_{k}^{\prime},\chi_{d})|$ are provided for every zero, with all bounds below $10^{-447}$. This dataset extends and supersedes the tables in Appendices B and C of the present paper.